### v3

# Tiny-Align: Bridging Automatic Speech Recognition and Large Language Model on Edge

Abstract
The combination of Large Language Models (LLM) and Automatic Speech Recognition (ASR), when deployed on edge devices (called edge ASR-LLM), can serve as a powerful personalized assistant to enable audio-based interaction for users. Compared to text-based interaction, edge ASR-LLM allows accessible and natural audio interactions. Unfortunately, existing ASR-LLM models are mainly trained in high-performance computing environments and produce substantial model weights, making them difficult to deploy on edge devices. More importantly, to better serve users’ personalized needs, the ASR-LLM must be able to learn from each distinct user, given that audio input often contains highly personalized characteristics that necessitate personalized on-device training. Since individually fine-tuning the ASR or LLM often leads to suboptimal results due to modality-specific limitations, end-to-end training ensures seamless integration of audio features and language understanding (cross-modal alignment), ultimately enabling a more personalized and efficient adaptation on edge devices. However, due to the complex training requirements and substantial computational demands of existing approaches, cross-modal alignment between ASR audio and LLM can be challenging on edge devices. In this work, we propose a resource-efficient cross-modal alignment framework that bridges ASR and LLMs on edge devices to handle personalized audio input. Our framework enables efficient ASR-LLM alignment on resource-constrained devices like Raspberry Pi 5 (8GB RAM), achieving 50x training time speedup while improving the alignment quality by more than 50%. To the best of our knowledge, this is the first work to study efficient ASR-LLM alignment on resource-constrained edge devices.
I Introduction
The Large Language Model (LLM) deployed on the edge device (edge LLM) can serve as a personalized assistant, guaranteeing data privacy [1] and continuous service without relying on stable internet connections [2, 3, 4]. Edge LLMs normally interact with users based on text (text-based interaction). Yet, it is more natural and accessible for users to interact with edge LLMs, their personalized assistants, via audio [5]. To enable the audio-based interaction, the Automatic Speech Recognition (ASR) models can be employed [6] to simply take the audio as the input and send the converted text into the edge LLM [7]. Existing works, on the other hand, have shown that this simple combination of ASR and LLM can fail in cases where the audio has no corresponding text [8], or the mismatching pre-trained knowledge in the ASR and the edge LLM can compromise the performance of their combination.
Figure 1: Evaluation of ASR-LLM alignment using Whisper-base and Gemma-2-2B with and without GPTQ (G) compression. Comparison includes three existing approaches (A1, A2, A3) and a preliminary design using only projector and LLM embedding (off-the-Shelf). This preliminary result motivates our work (detailed in Fig. 2), which outperforms existing approaches.
Different from simply combining the ASR and the edge LLM, past works [9] concatenate them at the feature level via the projector. In this concatenation, the ASR functions as the audio encoder and produces the audio embedding, which will then be sent to the edge LLM for LLM generation (edge ASR-LLM). Usually by default, the projector is a simple multi-layer perception (MLP) that matches the dimensionality of the audio embedding with the LLM. While the feature-level concatenation can overcome the problems in the simple combination of the ASR and the edge LLM, it often requires end-to-end training so the audio embeddings can be properly understood by the edge LLM (cross-modal alignment).
For cross-modal alignment, there are three main approaches to end-to-end training, as shown right part of Fig. 1. The first approach (A1) freezes both the ASR encoder and LLM while training only the projector [10]. This method updates the projector parameters by comparing LLM-generated outputs: one from projector-transformed audio embeddings and another from the corresponding text input. The second approach (A2) extends the first approach by jointly training [11] both the projector and LLM, creating a larger learning space. The third approach (A3) takes a different direction by training only the ASR encoder [12] while keeping the projector and LLM frozen.
The three approaches, while demonstrating decent performances in cross-modal alignment, can involve a huge size of parameter calculation upon the complex hyperparameters and the model tuning space. In this way, the huge resource budget can be prohibitive for cross-modal alignment on edge devices. Under the existing cross-modal alignment approaches, even if we choose the compressed model with smaller parameters, we can expect either significant performance downgrading due to the lowering of parameter size or the still prohibitively high resource usage due to the internal training process design. As shown left part of Fig. 1, we provide a pre-evaluation of the three alignment approaches on the concatenation of an ASR model–Whisper [13] and the edge LLM–Gemma-2-2B [14]. We also employ the compress Gemma-2-2B based on GPTQ [15] (G). The ASR and the edge LLM are connected by a two-layer MLP. In terms of training time (latency) and performance, we do not observe significant differences between the original model and its compressed version.
For approaches A1 and A2, the LLM generation involved can consume a huge amount of resources, leading to a longer convergence time. The approach of A2, which takes the projector and the LLM for training, outperforms the other two approaches but it also takes the most resources. The approach of A1 is more lightweight, but it still uses the LLM, and its reduction in finetuning resources leads to significant performance degradation. Meanwhile, The A1’s lower performance than the A2 can be due to the simple projector design. As shown in Fig. 1, while the A2 obtains decent performance, it also has the highest latency mostly due to the LLM generation. The off-the-shelf is tentatively made to avoid the LLM generation. To do that, we pick and use only its embedding layer to avoid the LLM generation. The embedding layer interfaces directly with the ASR encoder through a well-designed projector, which requires both sufficient embedding capacity and architectural compatibility with the ASR encoder and the rest components of the edge LLM. In Fig. 1, we compare this design as off-the-shelf with the three existing approaches. Surprisingly, we discovered that this design can outperform A1, A2, and even A3 in terms of performance and training latency.
Motivated by our observations, in this paper, we propose a resource-efficient alignment framework to enable cross-modal alignment between ASR and LLM on-edge devices, effectively balancing resource usage and end-to-end training performance. In our alignment framework, we design a novel projector that provides strong expressive power while consuming minimal training resources. Given its effectiveness in edge ASR-LLM scenarios, we name our framework as Tiny-Align. Our experiments demonstrate that Tiny-Align achieves significant performance improvements in aligning ASR encoders with various edge LLMs, requiring minimal time and computational resources.
Our major contributions can be summarized as follows:
  • We propose a novel approach (Tiny-Align) to enable the efficient alignment of ASR and LLM on resource-constrained edge devices. To the best of our knowledge, Tiny-Align is the first of its kind. Compared to existing alignment methods that require server-scale computing resources, Tiny-Align achieves training convergence 50$\times$ faster while Improving the quality of the results under ROUGE-1 and ROUGE-L scores by more than 50%.
  • Tiny-Align employs a new projector design (BridgeFormer) based on a transformer encoder architecture but without the positional encoding component. BridgeFormer provides a larger embedding space than existing MLPs or DNN-based designs and is shown to outperform the latter.
  • We further introduce a novel instruction injection mechanism for Tiny-Align’s inference, and show that it can further improve the quality of the results by about 50% compared to Tiny-Align without instruction injection.
All our results have been conducted on datasets collected from individuals with various disabilities, such as dementia, aphasia, and speech-language impairments (SLI), hence exhibiting a wide range of speech needs. We show that Tiny-Align can support high-quality interactions for them through an ASR interface.
Figure 2: Tiny-Align Framework. It unifies the trainable parameters into the designed transformer-based projector (BridgeFormer), which is trained to minimize the loss between the audio embedding made by itself and the textual embedding made through the original LLM embedding layer.
II Background
II-A Automatic Speech Recognition onto edge LLM
Extending edge LLM’s capability to process audio input can significantly enhance user accessibility. ASR models serve as the bridge between audio and text modalities, with several prominent approaches offering different trade-offs between performance and resource efficiency.
Wav2vec2 [16] pioneered an efficient approach to audio processing by directly learning representations from raw waveforms. Through a multi-layer convolutional network, it extracts context-dependent representations that effectively capture speech patterns. The model employs contrastive learning to identify relevant features and discern patterns in the audio signal. While its streamlined architecture makes it particularly efficient for edge deployment, wav2vec2 has demonstrated robust performance in various scenarios [17], including child speech recognition, highlighting its capability to effectively capture speech features across diverse conditions.
Whisper [13] and Conformer [18] represent another advancement in ASR technology, leveraging transformer architectures for improved accuracy. These models process audio through self-attention mechanisms to capture long-range dependencies in speech signals, achieving superior performance in complex transcription tasks across various languages and acoustic conditions [19]. However, their sophisticated architecture requires more computational resources compared to wav2vec2.
More recent approaches like Diffsound [20], AudioLM [21], and Tango2 [22] adopt generative frameworks that extend beyond simple transcription. Based on diffusion models or autoregressive generation, these ASRs can not only recognize speech but also model the underlying audio distribution. While these models achieve state-of-the-art performance in complex scenarios, their resource-intensive nature makes them highly challenging to deploy in edge environments without significant optimization.
II-B Cross-modal Alignment on Edge
One key challenge in personalizing edge LLM is adapting to users’ preferred input modality while operating within the constraints of limited computational resources [23, 24, 25, 26]. Cross-modal alignment enables the integration of non-textual inputs by combining specialized machine learning models for various modalities with the edge LLM which provides advanced reasoning capabilities for analyzing cross-modal data [27]. However, this alignment process demands substantial resources, making deployment on edge devices particularly challenging [28, 29]. For example, on resource-constrained devices like Raspberry Pi 5 (8GB RAM, quad-core Cortex-A76), cross-modal alignment can quickly exhaust available memory and computational resources, with processing latency becoming prohibitive for real-time applications. These constraints highlight the need for optimizing cross-modal alignment specifically for edge deployment.
II-C Edge ASR-LLM Alignment for People with Speech Difficulties
While edge ASR-LLM alignment enhances accessibility for general users, it holds particular promise for individuals with speech difficulties. In this work, we demonstrate the effectiveness of our alignment approach using datasets from individuals with dementia, aphasia, and specific language impairments (SLI), showcasing how edge ASR-LLM can adapt to and understand users despite significant speech challenges. This capability becomes increasingly crucial given the growing population affected by these conditions—over 6.7 million people with dementia, 2 million with aphasia, and more than 180,000 new cases annually [30]. For these individuals, traditional text-based interaction is infeasible due to motor and cognitive challenges, making voice-based interaction not merely an alternative, but a necessity for healthcare delivery.
III Proposed Work
In this section, we first provide an overview of our Tiny-Align framework shown in Fig. 2. We then delve into the technical details, beginning with the pivotal component: the projector design. Following this, we explain how the projector collaborates with the LLM embedding layer. We then elaborate on the types of ASR models suitable for our edge ASR-LLM alignment framework. Finally, we introduce our instruction injection method, which enables projector-generated embeddings to better facilitate LLM generation.
III-A Framework Overview
As shown in Fig. 2, Tiny-Align operates in two modes. During the training mode (upper half in Fig. 2), our proposed BridgeFormer establishes a shared embedding space through the proposed EmbedLink, enabling effective cross-modal alignment between ASR and LLM, guided by the LLM embedding layer. Additionally, we analyze and identify optimal ASR models that balance feature extraction capability with resource efficiency, ensuring Tiny-Align’s effectiveness across different deployment scenarios. During the inference model (lower half in Fig. 2), ASR-extracted features are transformed via BridgeFormer into LLM-compatible embeddings, which can concatenate with instructions under our proposed instruction injection mechanism. Either the LLM-compatible embeddings or the concatenations will go through the LLM for generating the output.
III-B BridgeFormer: Transformer-based Projector
Motivated by our preliminary study in Fig. 1, where a simple MLP projector demonstrates promising results in edge ASR-LLM alignment, we design an enhanced projector architecture that offers a larger audio embedding space where the embeddings can also be understood by the LLM, while maintaining resource efficiency. This design is driven by two key requirements: the ability to support continuous learning during user interactions and architectural compatibility with LLMs to ensure effective embedding transformation. As the sole trainable module in Tiny-Align, our projector aims to leverage the computational resources saved from avoiding LLM generation and training, allowing for a moderately larger architecture while remaining within edge device constraints.
To meet these design criteria, we introduce a transformer encoder architecture into the projector, named BridgeFormer to reflect its role in bridging ASR and LLM models. Compared to classical MLPs or deep neural networks, the transformer architecture offers several advantages: a larger embedding space with enhanced learning potential, natural alignment with LLM’s architecture (as transformers were originally designed for natural language processing) [31, 32], and proven effectiveness in handling sequential data. Furthermore, since BridgeFormer processes ASR-extracted audio features where temporal positioning is already encoded, we remove the traditional positional encoding component, simplifying the architecture and improving its compatibility with both ASR and edge LLM.
Figure 3: Architecture of BridgeFormer.
Consider audio features extracted from an ASR encoder with shape $[1,N,D_{a}]$, where $N$ represents the audio sequence length and $D_{a}$ is the ASR embedding dimension. BridgeFormer first employs an input MLP to project these features into a higher-dimensional space, creating initial embeddings of shape $[1,N,H]$, where $H$ is the hidden dimension. These embeddings then pass through the transformer encoder, where $M$ attention heads across $E$ encoder layers to process and refine the features while maintaining the sequence dimensions. To prevent overfitting and control sequence length, an adaptive pooling layer reshapes the sequence to a predefined token size $T$, producing outputs of shape $[1,T,H]$. Finally, an output MLP transforms these features to match the LLM embedding dimension $D_{l}$, resulting in LLM-compatible embeddings of shape $[1,T,D_{l}]$. This process can be formally expressed as:
$E_{\text{in}}=\text{MLP}_{\text{input}}(F_{\text{audio}})\in\mathbb{R}^{1\times N\times H}$ (1)
where $F_{\text{audio}}\in\mathbb{R}^{1\times N\times D_{a}}$ represents the input audio features from ASR.
$E_{\text{trans}}=\text{Transformer}_{M,E}(E_{\text{in}})\in\mathbb{R}^{1\times N\times H}$ (2)
where $M$ denotes the number of attention heads and $E$ represents the number of encoder layers.
$E_{\text{pool}}=\text{Pool}(E_{\text{trans}})\in\mathbb{R}^{1\times T\times H}$ (3)
where $T$ is the target token size (discussed in the following section).
$E_{\text{out}}=\text{MLP}_{\text{output}}(E_{\text{pool}})\in\mathbb{R}^{1\times T\times D_{l}}$ (4)
where $D_{l}$ represents the LLM embedding dimension.
BridgeFormer’s architecture is designed for flexibility and resource adaptability. The model can be easily scaled to match available edge resources through three key parameters: hidden dimension (controlling embedding capacity), number of attention heads (affecting feature relationship modeling), and number of transformer encoder layers (determining processing depth). As shown in Fig. 3, BridgeFormer comprises four essential components: (1) an input MLP that projects features into a high-dimensional embedding space, (2) transformer encoder layers that process and refine these embeddings, (3) a pooling layer that prevents overfitting and controls dimensionality, and (4) an output MLP that generates the final LLM-compatible embeddings.
III-C EmbedLink: Cross-modal Alignment via Embedding Layer
With BridgeFormer’s architecture established, we now focus on its training methodology for creating a shared embedding space. We propose EmbedLink, a simple yet effective pipeline that avoids resource-intensive LLM generation during cross-modal alignment. Our approach leverages a key insight: LLM’s text processing fundamentally relies on embeddings from its embedding layer. Therefore, if BridgeFormer can transform audio features into embeddings that closely match those generated by the LLM’s embedding layer from corresponding text input, we can achieve effective cross-modal alignment without the computational overhead of full LLM inference.
Figure 4: Demonstration of the EmbedLink usage in the BridgeFormer training. The EmbedLink uses casted token size, LLM embedding layer, and two loss functions to update parameters in BridageFormer.
A critical challenge in this alignment process is managing the dimensional disparities between audio features and LLM embeddings. As illustrated in Fig. 4, audio embeddings have dimensions determined by audio length $N$ and ASR encoding dimension $D_{a}$, while LLM embeddings are shaped by token size and embedding dimension $D_{l}$. The audio length $N$ typically exceeds the token count significantly due to pauses, silence periods, and other audio-specific characteristics. Directly matching these dimensions could either result in learning failure or significant information loss.
To address this dimensional mismatch, EmbedLink introduces a casted token size $T$, which serves as a fixed intermediate dimension between typical text token counts and audio lengths. After system initialization, this size remains constant to ensure training stability. For text inputs exceeding $T$ tokens, EmbedLink performs truncation; for shorter inputs, it applies padding. Meanwhile, BridgeFormer’s output dimension is set to match the LLM embedding dimension $D_{l}$ (typically 2048 for models like Llama).
Consider an audio input processed by ASR into features $F_{\text{audio}}$ of shape $[1,N,D_{a}]$ and its corresponding text. The LLM embedding layer first processes the text input: if its token count exceeds the casted token size $T$, it truncates the excess; if shorter, it applies padding to reach $T$. This processed text is then converted by the LLM embedding layer into embeddings $E_{\text{text}}$ of shape $[1,T,D_{l}]$. Meanwhile, BridgeFormer transforms $F_{\text{audio}}$ through its architecture (as described in the previous section) to produce $E_{\text{audio}}$ with matching dimensions $[1,T,D_{l}]$. This alignment process can be formally expressed as:
$E_{\text{text}}=\text{LLM}_{\text{embed}}(X)\in\mathbb{R}^{1\times T\times D_{l}}$ (5)
where $X$ represents the input text after token size adjustment.
$E_{\text{audio}}=\text{BridgeFormer}(F_{\text{audio}})\in\mathbb{R}^{1\times N\times D_{a}}\rightarrow\mathbb{R}^{1\times T\times D_{l}}$ (6)
To optimize BridgeFormer’s parameters, we employ a combined loss function that considers both direct feature matching and semantic similarity:
$\mathcal{L}=\alpha\cdot\text{MSE}(E_{\text{audio}},E_{\text{text}})+\beta\cdot(1-\text{cos}(E_{\text{audio}},E_{\text{text}}))$ (7)
where $\alpha$ and $\beta$ are weighting coefficients balancing the two objectives.
III-D Feature-based ASR encoder
After establishing BridgeFormer and EmbedLink, selecting an appropriate ASR encoder becomes crucial for our framework. While we adopt existing ASR models rather than designing new ones, the selection must balance hardware efficiency with feature extraction capability. As discussed in Section II-A, we consider three mainstream ASR types: feature-based ASR (e.g., wav2vec2-base), transformer-based ASR (e.g., Whisper-base), and generative ASR (e.g., AudioLDM).
Our analysis combines theoretical representation space study with empirical evaluation of resource usage and training efficiency. For theoretical analysis, we examine how these ASR models map an input audio sequence $x\in\mathcal{X}$ to their respective feature spaces:
$F_{\text{audio}}^{\text{gen}}=\text{ASR}_{\text{gen}}(x)\in\mathbb{R}^{1\times N\times D_{g}}$ (8)
where generative ASRs like AudioLDM produce variable-dimension features $D_{g}$ adapted to input complexity, potentially capturing the most comprehensive audio information but requiring substantial computational resources.
$F_{\text{audio}}^{\text{feat}}=\text{ASR}_{\text{feat}}(x)\in\mathbb{R}^{1\times l(x)\times D_{f}}$ (9)
where feature-based ASRs like wav2vec2 maintain length proportional to input ($l(x)$) with fixed feature dimension $D_{f}$ (typically 768), offering a balanced representation of audio features.
$F_{\text{audio}}^{\text{trans}}=\text{ASR}_{\text{trans}}(x)\in\mathbb{R}^{1\times L\times D_{t}}$ (10)
where transformer-based ASRs like Whisper use fixed dimensions ($L=1500,D_{t}=512$), providing consistent but potentially constrained feature representation.
Our empirical evaluation, as shown in Fig. 5, reveals two critical insights about hardware efficiency and feature quality. First, examining DRAM usage during feature extraction, we observe that generative ASRs consume significantly more memory (over 3500MB) compared to feature-based (458MB) and transformer-based (322MB) alternatives. This substantial memory footprint could severely limit the resources available for parameter learning on edge devices. Second, we evaluate feature quality through a training efficiency test using a naive MLP, tracking the cross-entropy loss over 1000 epochs. While generative ASR features show gradual, smooth convergence, feature-based ASR demonstrates remarkably rapid convergence within 400 epochs, with wav2vec2 achieving near-zero loss. Transformer-based ASR shows similar quick convergence but with slightly higher final loss values.
This comprehensive analysis leads us to select feature-based ASR, particularly wav2vec2, as the optimal encoder for our Tiny-Align framework. It offers the best balance between hardware efficiency and feature extraction capability, crucial for effective edge deployment while maintaining high-quality cross-modal alignment.
Figure 5: Resource efficiency comparison of three ASR models. Left: Peak RAM usage during audio feature extraction. Right: Training convergence curves using a naive MLP with features extracted from different ASR models.
III-E Instruction Injection
When the proper ASR encoder is chosen and EmbedLink has trained BridgeFormer to generate LLM-compatible embeddings, these embeddings purely represent the audio input information. However, to elicit high-quality LLM responses, instructions are often required [33], and these may vary across different use cases. Rather than incorporating instructions during BridgeFormer training, we design a flexible instruction injection module for inference time, allowing users to adapt instructions based on specific needs. As illustrated by the deep blue dashed lines in Fig. 2, this module implements a streamlined instruction injection process.
Given a task-specific instruction $I$, we first convert it into LLM-recognizable embeddings:
$E_{\text{inst}}=\text{LLM}_{\text{embed}}(I)\in\mathbb{R}^{1\times N_{i}\times D_{l}}$ (11)
where $N_{i}$ represents the instruction sequence length and $D_{l}$ is the LLM embedding dimension. These instruction embeddings are then concatenated with the BridgeFormer output along the sequence dimension:
$E_{\text{final}}=\text{Concat}(E_{\text{inst}},E_{\text{audio}})\in\mathbb{R}^{1\times(N_{i}+T)\times D_{l}}$ (12)
where $T$ is the casted token size used in BridgeFormer.
This concatenated embedding sequence forms a complete input for the LLM, where the instruction embeddings guide the model’s interpretation of the subsequent audio-derived embeddings. This modular design allows for flexible instruction modification without requiring retraining of the alignment model, making it particularly suitable for edge deployment where adaptation to different tasks may be necessary.
IV Experimental Evaluation
IV-A Experimental Setup
IV-A1 Datasets
To demonstrate our Tiny-Align framework, we employ five diverse datasets from TalkBank [34], a comprehensive database for language research. These datasets include: (1) ADReSS-IS2020 [35], containing speech recordings of Alzheimer’s patients describing the Cookie Theft picture; (2) Baycrest [36], featuring structured interviews with dementia patients; (3) EllisWeismer [37], focusing on children with specific language impairments; (4) ENNI [38], containing narrative samples from children with language disorders; and (5) NEURAL [39], comprising speech samples from individuals with various neurological conditions. Each dataset contains recordings from over twenty patients, with paired audio samples and corresponding transcripts. For training and evaluation, we segment these recordings into sentence-level audio-transcript pairs, resulting in approximately X training pairs and Y validation pairs. Our usage of these privacy-concerned datasets is approved by IRB [40].
IV-A2 LLM Models
We evaluate our framework across multiple state-of-the-art LLMs, including Llama-3.2-1B [41], Llama-3.2-3B [41], Phi-3.5-mini [42], Gemma-2-2B [14], and StableLM-2-1.6B [43]. These models represent different architectures and parameter scales, allowing a comprehensive evaluation of our alignment approach. For LLM configuration, we employ maximum length padding in the tokenizer and set the temperature to 0.1 with top-k sampling (k=50) to ensure consistent and high-quality outputs.
IV-A3 Default Experimental Setting
For audio feature extraction, we utilize the wav2vec2-base-960h model [16] as our ASR model. It was pre-trained on 960 hours of LibriSpeech data. All audio inputs are preprocessed with a sampling rate of 16 kHz to match the model’s training conditions. Our transformer-based projector employs a compact yet effective architecture with 4 attention heads, a hidden dimension of 256, and 4 transformer layers. We empirically set the casted token size to 30, which adequately captures typical sentence lengths while maintaining computational efficiency.
The training process uses the AdamW optimizer with an initial learning rate of 1e-3 and a linear decay schedule. This configuration balances training stability with convergence speed, which is particularly important for edge deployment scenarios where training resources may be limited.
The preliminary validation on a Rasberry Pi 5 can be found in section IV-B1, followed by the comprehensive experiments run on a single Nvidia P100 GPU.
IV-A4 Evaluation Methods
We evaluate our framework from both effectiveness and efficiency perspective. For effectiveness, we employ a dual-path comparison approach: given an audio input, we compare the output from our ASR-LLM system (OutputA) with the output generated by directly feeding the ground truth transcript to the LLM (OutputL). The similarity between these outputs is measured using ROUGE scores [44], including ROUGE-1 and ROUGE-L, which assess word-level and longest common subsequence matches, respectively. For the ROUGE scores, higher values represent better performance of the LLM generated content.
For efficiency evaluation, we focus on convergence time as our primary metric, which measures not only the training speed but also resource utilization efficiency [45]. Extended training periods on edge devices can lead to sustained high DRAM usage, GPU occupancy, and power consumption. We define convergence as the point where the loss stabilizes ($\Delta\text{loss}<\varepsilon$), analyzing both time-to-convergence and total resource consumption during training. This metric is particularly crucial for edge deployment, where prolonged training can strain limited device resources and impact system availability. To evaluate personalization effectiveness, we conduct user-level performance analysis by randomly sampling 20 users from each dataset and computing their average performance metrics.
IV-A5 Baselines
To evaluate our Tiny-Align framework, we compare it with several state-of-the-art cloud-based models. While these baselines leverage extensive computational resources and large-scale pretraining to achieve high performance, our focus remains on efficient edge deployment through training our designed projector. We select three representative baselines for comparison, where each baseline can correspond to one approach shown in Fig. 1. NExT-GPT [10], which incorporates LLM generation during ASR-LLM alignment and computes loss based on the generated content, can represent A1.X-VILA [11], which simultaneously fine-tunes both the connecting projector and the LLM during alignment, can represent A2. LanguageBind [12], which trains the ASR encoder in the edge ASR-LLM system while freezing the projector and the LLM, can represent A3. The A1, A2, A3 refer to the approaches described in Secion I. Compared to our method, these baselines involve more extensive parameter calculations and require substantially more computational resources.
Figure 6: Demonstration of the viability of Tiny-Align on a resource-constrained edge device (Raspberry Pi 5), evaluated under various configurations (x-axis) of the proposed BridgeFormer. The default settings are marked in dark green.
IV-B Results
IV-B1 Edge Device Implementation Validation
To assess whether edge device can support running the full model pipeline (i.e., ASR, projector, and LLM), we perform a preliminary validation on a Raspberry Pi 5. In this experiment, we evaluate two critical factors: throughput and RAM usage.
As shown in Figure 6, we explore various configurations of our BridgeFormer. The left panel illustrates the trends in throughput and memory usage across different hidden dimensions, while the right panel shows the impact of varying the number of attention heads.
Figure 7: Demonstration of the loss decrease when training the wav2vec2+Gemma-2-2B based on the four methods on two datasets.
IV-B2 Comprehensive Experiments
In this set of the experiments, we first analyze the training loss convergence of our method compared to three baselines. We evaluate an ASR-LLM system composed of wav2vec2 and Gemma-2-2B on two datasets: ADReSS-IS2020 and ENNI. Training loss is tracked over 1000 epochs. As shown in Fig. 7, our method achieves convergence within 400 epochs on ADReSS-IS2020 and demonstrates even faster convergence (within 100 epochs) on the ENNI dataset. This rapid convergence enables efficient ASR-LLM alignment while minimizing computational resource usage and training time. Additionally, we can observe the drastic fluctuations of the loss curve for the three baselines, indicating the training under these three methods may be less stable. One thing to note is that, since the baselines are designed without considering the resource restrictions, we must limit their trainable parameters so that they can run in resource-limited environments.
Figure 8: Evaluation of our method and baselines on different training data sizes given ADReSS dataset. We choose wav2vec2 and Llama-3.2-1B to formalize the ASR-LLM system. Convergence Time and ROUGE-1 score are used to measure the training time (latency) and performance.
In addition to the training loss, data size is a crucial factor in evaluating edge ASR-LLM alignment. An ideal alignment method should maintain both efficiency and effectiveness across varying data sizes, ensuring compatibility with edge devices even when users possess large volumes of data. To evaluate this aspect, we conduct experiments using the ADReSS-IS2020 dataset, creating five training data groups of different sizes. We analyze both Convergence Time and ROUGE-1 scores using our method and three baselines on an ASR-LLM system composed of wav2vec2 and Llama-3.2-1B. As shown in Fig. 8, our method exhibits significantly shorter convergence times and demonstrates better scaling behavior across all data sizes compared to the baselines. Furthermore, our method maintains effective ASR-LLM alignment even with limited resources, achieving comparable performance ratios between small and large data sizes. This consistent performance across different data scales demonstrates the robustness and efficiency of our method regardless of data volume.
TABLE I: Evaluation of Tiny-Align (Ours), NExT-GPT (A1), X-VILA (A2), and LanguageBind (A3) on five datasets, when they are used to align the ASR and five LLM alignment. During the performance evaluation, ROUGE-1 (R-1) and ROUGE-L (R-L) are involved as the metrics and limit the training epoch up to 400. During the Latency evaluation, convergence time (C-T) is used and does not limit the training epoch to 400.
Dataset Method ASR + Llama-3.2-1B ASR + Llama-3.2-3B ASR + Gemma-2-2B ASR + Phi-3.5-mini ASR + StableLM-2-1.6B
R-1 R-L C-T(10s) R-1 R-L C-T(10s) R-1 R-L C-T(s) R-1 R-L C-T(10s) R-1 R-L C-T(10s)
ADReSS A1 0.021 0.025 293 0.032 0.027 293 0.051 0.049 2249 0.103 0.072 2520 0.030 0.030 1539
A2 0.152 0.138 3107 0.134 0.128 3107 0.196 0.112 7338 0.213 0.119 11875 0.093 0.091 3670
A3 0.104 0.096 1128 0.108 0.103 1128 0.185 0.119 1841 0.029 0.026 2093 0.064 0.060 1577
Ours 0.192 0.164 111 0.205 0.193 111 0.268 0.154 97 0.225 0.121 104 0.180 0.176 93
Baycrest A1 0.024 0.024 792 0.024 0.024 792 0.017 0.016 11517 0.050 0.047 7375 0.003 0.003 5055
A2 0.141 0.104 949 0.152 0.109 949 0.183 0.114 4614 0.202 0.110 3847 0.095 0.093 5357
A3 0.065 0.057 3015 0.071 0.067 3015 0.088 0.060 9951 0.012 0.011 6903 0.019 0.011 4918
Ours 0.184 0.167 167 0.197 0.173 167 0.233 0.141 133 0.205 0.112 183 0.106 0.103 159
Ellis A1 0.025 0.032 873 0.032 0.081 873 0.062 0.059 10593 0.009 0.009 5372 0.019 0.016 6668
A2 0.187 0.145 1182 0.190 0.104 1182 0.193 0.107 3257 0.052 0.033 3526 0.044 0.044 2549
A3 0.035 0.027 1563 0.045 0.051 1209 0.014 0.015 2371 0.022 0.010 6784 0.017 0.013 5366
Ours 0.223 0.221 95 0.223 0.221 95 0.196 0.135 91 0.146 0.079 136 0.091 0.091 153
ENNI A1 0.011 0.009 503 0.015 0.001 503 0.034 0.033 12530 0.109 0.076 13002 0.109 0.031 3507
A2 0.109 0.089 742 0.157 0.131 742 0.191 0.115 2770 0.219 0.128 2165 0.164 0.164 1024
A3 0.034 0.034 1708 0.034 0.034 1708 0.256 0.157 5407 0.085 0.073 5461 0.015 0.009 2141
Ours 0.124 0.103 91 0.143 0.135 91 0.244 0.151 82 0.214 0.128 111 0.303 0.268 74
NEURAL A1 0.023 0.023 1506 0.023 0.023 1506 0.054 0.048 17700 0.116 0.088 9760 0.009 0.009 9097
A2 0.075 0.067 2187 0.075 0.067 2187 0.177 0.108 2645 0.157 0.087 3875 0.060 0.057 2356
A3 0.051 0.048 1709 0.021 0.018 1683 0.034 0.045 3315 0.105 0.096 6149 0.061 0.055 6533
Ours 0.104 0.098 141 0.158 0.109 141 0.219 0.132 144 0.172 0.093 217 0.129 0.124 184
Following our analysis of data size impact, we conduct a comprehensive evaluation of our method against the baselines. Based on our training loss analysis, we set the training epoch limit to 400 to reflect resource-constrained scenarios in ASR-LLM alignment. Given that our previous study did not show significant performance variations for different data sizes, we use a dataset of 100 samples for this comprehensive evaluation. As shown in TABLE I, our proposed method, Tiny-Align, demonstrates superior performance across five datasets and five ASR-LLM systems, achieving significantly higher ROUGE-1 and ROUGE-L scores compared to the three baselines. Furthermore, we analyze convergence time without the 400-epoch constraint, setting a target loss of 0.05. The results, presented in the third column for each ASR-LLM system in TABLE I, show that Tiny-Align achieves significantly faster convergence compared to the baseline methods.
Figure 9: Performance comparison of instruction injection in our Tiny-Align framework across five datasets and two ASR-LLMs. The training process is unrestricted by epoch limits, allowing full convergence.
Furthermore, we evaluate the instruction injection mechanism on two ASR-LLM combinations: wav2vec2+Gemma-2-2B and wav2vec2+StableLM-2-1.6B. In these experiments, we remove the epoch limit to allow complete BridgeFormer training. Instructions are then injected during the inference stage. As shown in Fig. 9, instruction injection significantly improves LLM generation performance with the same trained BridgeFormer, indicating enhanced LLM comprehension of BridgeFormer-produced audio embeddings.
V Conclusion
In this paper, we present the Tiny-Align framework, which enables edge ASR-LLM alignment. By properly designing a transformer-based projector–BridgeForme–and training it under EmbedLink, Tiny-Align can significantly lower the training cost. BridgeFormer can both effectively and efficiently build a shared embedding space, where the audio embeddings can be properly converted into LLM-compatible embeddings. With the proper ASR encoder chosen, and the boost from instruction injection, Tiny-Align results in significantly higher performance in edge ASR-LLM alignment.